# Video-MME: The First-Ever Comprehensive Evaluation Benchmark of Multi-modal LLMs in Video Analysis

### v2

@@ Abstract
In the quest for artificial general intelligence, Multi-modal Large Language Models (MLLMs) have emerged as a focal point in recent advancements. However, the predominant focus remains on developing their capabilities in static image understanding. The potential of MLLMs in processing sequential visual data is still insufficiently explored, highlighting the absence of a comprehensive, high-quality assessment of their performance. In this paper, we introduce Video-MME, the first-ever full-spectrum, Multi-Modal Evaluation benchmark of MLLMs in Video analysis. Our work distinguishes from existing benchmarks through four key features: 1) Diversity in video types, spanning 6 primary visual domains with 30 subfields to ensure broad scenario generalizability; 2) Duration in temporal dimension, encompassing both short-, medium-, and long-term videos, ranging from 11 seconds to 1 hour, for robust contextual dynamics; 3) Breadth in data modalities, integrating multi-modal inputs besides video frames, including subtitles and audios, to unveil the all-round capabilities of MLLMs; 4) Quality in annotations, utilizing rigorous manual labeling by expert annotators to facilitate precise and reliable model assessment. 900 videos with a total of 254 hours are manually selected and annotated by repeatedly viewing all the video content, resulting in 2,700 question-answer pairs. With Video-MME, we extensively evaluate various state-of-the-art MLLMs, including GPT-4 series and Gemini 1.5 Pro, as well as open-source image models like InternVL-Chat-V1.5 and video models like LLaVA-NeXT-Video. Our experiments reveal that Gemini 1.5 Pro is the best-performing commercial model, significantly outperforming the open-source models with an average accuracy of 75%, compared to 71.9% for GPT-4o. The results also demonstrate that Video-MME is a universal benchmark, which applies to both image and video MLLMs. Further analysis indicates that subtitle and audio information could significantly enhance video understanding. Besides, a decline in MLLM performance is observed as video duration increases for all models. Our dataset along with these findings underscores the need for further improvements in handling longer sequences and multi-modal data, shedding light on future MLLM development. Project page: https://video-mme.github.io.
@@ 1 Introduction
The rapid development of multi-modal large language models (MLLMs) in recent years [66, 48, 15, 50, 3, 72] has highlighted their impressive perception and cognitive capabilities across various multimodal benchmarks [13, 67, 42, 73]. These advancements demonstrate the great potential of MLLMs to serve as a foundation that can digest the multi-modal real world [34] and pave the way toward artificial general intelligence. However, current MLLMs and their evaluation primarily focus on static visual data understanding, which fails to capture the dynamic nature of the real world involving complex interactions between objects over time. To approximate real-world scenarios more accurately, it is crucial to explore and assess the capabilities of MLLMs on sequential visual data, such as videos. Many early efforts [71, 53, 24, 30] have been made to inspire the video understanding potentials of MLLMs with promising results. However, existing video-based benchmarks [29, 24, 45, 39] are still limited to thoroughly reveal their performance, such as a lack of diversity in video types, insufficient coverage of temporal dynamics, and the narrow focus on a single modality. These inevitably hinder the all-around evaluation of MLLMs.
Figure 1: Examples of Video-MME. The ground-truth answer is highlighted in green. In Video-MME, all data, including question-answering annotations, videos, subtitles, and audios, are manually collected and curated, ensuring diversity and quality.
@@
To this end, we introduce Video-MME, the first-ever comprehensive Multi-Modal Evaluation benchmark crafted for MLLMs in Video analysis. As exemplified in Figure 1, we meticulously curate a dataset of 900 videos across various scenarios, and annotate a set of 2,700 high-quality multiple-choice questions (3 per video) to foster a robust evaluation. As presented in Figure 2, for generalizability, our dataset widely spans 6 visual domains, including Knowledge, Film & Television, Sports Competition, Artistic Performance, Life Record, and Multilingual, with 30 fine-grained categories, e.g., astronomy, technology, documentary, news report, esports, magic show, and fashion. Importantly, the videos vary significantly in length, ranging from 11 seconds to 1 hour, specifically evaluating the adaptability of MLLMs across varying temporal contexts. Furthermore, Video-MME enriches the assessment by incorporating the associated subtitles and audio tracks, thereby enhancing the analysis of multi-modal inputs for video understanding.
@@
Using Video-MME, we benchmark various state-of-the-art MLLMs, including GPT-4V [48], GPT-4o [49], and Gemini 1.5 Pro [54], alongside open-source image models like InternVL-Chat-V1.5 [9] and video models like LLaVA-NeXT-Video [74]. Our experiments in Table 4 indicate that Gemini 1.5 Pro is the highest-performing commercial model, achieving an average accuracy of 75%. In comparison, open-source MLLMs exhibit substantial gaps compared to commercial models. For instance, the leading open-source model, VILA-1.5 [31], attains an overall accuracy of 59%. These findings suggest there is considerable room for improvement in the open-source community. Our benchmark is also available to advanced image-based models by extending their input to multi-frame images, e.g., Qwen-VL-Max [5] and InternVL-Chat-V1.5 [9]. The accuracies of both the models reach 50% , which is close to that of the video specific model LLaVA-NeXT-Video, indicating that image understanding is the basis of video understanding, and the wide applicability of Video-MME in the filed of MLLMs. Further observations in Table 5 indicate that integrating subtitles and audios significantly enhances video comprehension capabilities, e.g., boosting Gemini 1.5 Pro by 6.2% and 4.3% respectively, with the gains being more pronounced for longer videos. A fine-grained analysis of task types reveals that subtitles and audios are particularly beneficial for videos requiring substantial domain knowledge. We also note a general decline in MLLM performance with increasing video length. This trend suggests that limitations in processing longer video sequences could be a critical bottleneck in the performance of MLLMs.
@@ 2 Video-MME
2.1 Dataset Construction
The dataset construction process of Video-MME consists of three steps: video collection, question-answering annotation, and quality review. The details are as follows.
@@ 2.1 Dataset Construction
Figure 2: (Left) Video categories. Our benchmark covers 6 key domains and 30 sub-class video types. (Right) Video duration length and question type distributions. Video-MME has a full spectrum of video length and covers different core abilities of MLLMs.
Video Collection.
For a comprehensive coverage of different video types, we first create a domain hierarchy for collecting raw videos from YouTube. We define 6 key domains: Knowledge, Film & Television, Sports Competition, Life Record, and Multilingual, based on popular tendencies on YouTube. Each domain is further divided into detailed tags, such as football and basketball for sports competition, resulting in a total of 30 fine-grained video classes. The full domain-tag hierarchy and its distribution can be found in the left part of Figure 2. For each class, we collect videos with varying duration lengths, including short (< 2 minutes), medium (4-15 minutes), and long videos (30-60 minutes). Besides, we also obtain corresponding meta-information such as subtitles (if provided) and audios for further investigation. Our final dataset consists of 900 videos with 744 subtitles and 900 audios, spanning various domains with relatively balanced duration lengths, as depicted in the right part of Figure 2.
Question-Answer Annotation.
After gathering the raw video data, we annotate it with high-quality question-answer (QA) pairs to evaluate the proficiency of MLLMs in interpreting video content. We employ a multiple-choice QA format to facilitate a straightforward and flexible assessment. The researchers proficient in English with extensive research experience in vision-language learning, perform the annotations. Specifically, they are first asked to watch the whole content of the video, and then to develop 3 corresponding questions, each with 4 potential options by repeatedly watching the video, contributing to 2,700 QA pairs in total. As shown in the bottom right corner of Figure 2, there are a total of 12 task types in the questions, including both perception, reasoning, and information synopsis. Particularly, each QA pair is required to be associated with the video content, avoiding MLLMs being able to answer without looking at the video.
Quality Review.
To guarantee the quality of our dataset, we conduct a rigorous manual review process. First, a different annotator is assigned to examine each QA pair to ensure that (i) language expressions are correct and unambiguous; (ii) the question is answerable, and the candidate options and provided golden option are reasonable. Furthermore, to ensure that the questions are challenging enough and require video content as a necessary condition [73], we provide the text-only questions to Gemini 1.5 Pro and filter out QA pairs that can be answered solely based on the textual questions. For example, the question “What is the biggest achievement of 10 of Argentina in 2022?” that can be directly inferred to the World Cup winner will be filtered out during this process. Questions that do not meet this criterion are returned to the original annotators for revision. By statistics, the accuracy of Gemini 1.5 Pro in the question-only setting is less than 15%. Through our rigorous dataset construction process, we strive to deliver a high-quality, diverse, and well-balanced dataset that will be instrumental for researchers in the field of multi-modal understanding.
2.2 Dataset Statistics
Here, we present the detailed statistics of our dataset to provide a more comprehensive understanding, including the meta information, QA pairs, certificate lengths, qualitative analysis, and comparison to previous works.
Video & Meta Information.
Our dataset comprises a total of 900 videos, 744 subtitles, and 900 audio files. Most videos are accompanied by both subtitles and audios, providing valuable resources for investigating the impact of external information on video understanding performance. The upper right part of Figure 2 illustrates the duration distribution of the collected videos. Specifically, within the short video category, longer videos occupy a larger portion. For medium-length videos, the duration is more uniformly distributed. In the category of long videos, there is a long-tailed distribution where longer videos have fewer samples. The bottom right part of Figure 2 shows the distribution of task types. Shorter videos predominantly involve perception-related tasks such as action and object recognition. In contrast, longer videos mainly feature tasks related to temporal reasoning. Overall, this analysis highlights that our dataset covers a wide range of video durations and various task types, enabling a comprehensive evaluation of temporal understanding.
QA Pair.
We demonstrate the language diversity of the questions and answers in our dataset. Table 2.2 lists the average word count of the textual fields in our dataset. The word counts for questions, options, and answers display notable consistency across different video lengths, suggesting a uniform style of QA pairs in our dataset. On the other hand, the word count for subtitles increases significantly with the length of the videos, e.g., short videos have an average word count of 198.6, while the long video subset have the count up to 6.5K. This trend indicates that longer videos contain more information, as evidenced by the increased volume of subtitles. The analysis reveals that our questions are diverse, and the answers are well-balanced. In addition, the distribution of the four answer options (A/B/C/D) follows a near-uniform distribution (25.1%/27.2%/25.3%/22.4%), ensuring an unbiased evaluation.
@@ 2.2 Dataset Statistics
Table 2: Average word count of different textual fields in Video-MME.
Dataset Question Options Answer Subtitles
Short 11.5 17.2 4.0 198.6
Medium 12.2 20.6 5.0 1425.6
Long 14.5 31.0 7.5 6515.6
All 12.7 22.9 5.5 3086.5
@@
Inspired by EgoSchema [45], we adopt the certificate length to analyze the temporal difficulty of the QA pairs. The certificate of a given video QA pair is defined as the minimum set of sub-clips of the video that are both necessary and sufficient to convince a human verifier that the marked annotation is correct. The certificate length is calculated as the sum of the temporal lengths of the sub-clips identified. We randomly sample 3 videos from each class and calculate the certificate length distribution with extra annotators. As shown in Table 1, our dataset yields a median certificate length of 26s, 164.7s, and 890.7s for short, medium and long videos, respectively. Compared with the certificate length of EgoSchema, our medium and long video subset requires much longer video content digestion to answer the question. To the best of our knowledge, this analysis makes our Video-MME the most challenging Video QA dataset to date.
@@
Comparison with Previous Benchmarks.
We compare the key difference of our dataset with previous benchmarks in Table 3. The first block lists traditional video benchmarks, which typically focus on specific domains such as TV videos or lack a clear hierarchy, making them less suitable for comprehensively diagnosing the limitations of MLLMs. In the middle block, although several benchmarks like TempCompass [39] and MVBench [24] explore multi-level evaluation and source videos from open domains, they still only cover videos with shorter durations. For example, the longest dataset, EgoSchema [45], includes videos up to 180 seconds, leaving the understanding of longer videos unaddressed. Our Video-MME is the first manually annotated benchmark that encompasses open-domain videos with durations ranging from 11 seconds to 1 hour. It evaluates different levels of video understanding ability and is supplemented with meta information such as subtitles and audios. This comprehensive approach uniquely positions Video-MME to advance the evaluation and development of MLLMs.

### v3

@@ Abstract
In the quest for artificial general intelligence, Multi-modal Large Language Models (MLLMs) have emerged as a focal point in recent advancements. However, the predominant focus remains on developing their capabilities in static image understanding. The potential of MLLMs to process sequential visual data is still insufficiently explored, highlighting the lack of a comprehensive, high-quality assessment of their performance. In this paper, we introduce Video-MME, the first-ever full-spectrum, Multi-Modal Evaluation benchmark of MLLMs in Video analysis. Our work distinguishes from existing benchmarks through four key features: 1) Diversity in video types, spanning 6 primary visual domains with 30 subfields to ensure broad scenario generalizability; 2) Duration in temporal dimension, encompassing both short-, medium-, and long-term videos, ranging from 11 seconds to 1 hour, for robust contextual dynamics; 3) Breadth in data modalities, integrating multi-modal inputs besides video frames, including subtitles and audios, to unveil the all-round capabilities of MLLMs; 4) Quality in annotations, utilizing rigorous manual labeling by expert annotators to facilitate precise and reliable model assessment. With Video-MME, we extensively evaluate various state-of-the-art MLLMs, and reveal that Gemini 1.5 Pro is the best-performing commercial model, significantly outperforming the open-source models with an average accuracy of 75%, compared to 71.9% for GPT-4o. The results also demonstrate that Video-MME is a universal benchmark that applies to both image and video MLLMs. Further analysis indicates that subtitle and audio information could significantly enhance video understanding. Besides, a decline in MLLM performance is observed as video duration increases for all models. Our dataset along with these findings underscores the need for further improvements in handling longer sequences and multi-modal data, shedding light on future MLLM development. Project page: https://video-mme.github.io.
@@ 1 Introduction
Figure 1: We introduce Video-MME to provide high-quality assessment of MLLMs’ performance, where all the videos and annotations are manually collected and curated. These are two highlighted examples of Video-MME with the ground-truth answers in green, which require the model to perform long-term contextual understanding and complex spatial and temperal reasoning. Video-MME is meticulously designed to pose significant challenges, thereby effectively evaluating the compositional video understanding capabilities of MLLMs.
Table 1: The comparison of various benchmarks encompasses several key aspects: the total number of videos (#Videos), the number of clips (#Clips), the average duration of the videos (Len.), the number of QA pairs (#QA Pairs), the method of annotation (Anno., M/A means the manually/automatic manner), the average number of QA pair tokens (QA Tokens), the average number of subtitle tokens (Sub. Tokens), whether the videos cover multiple duration levels (Multi-level), whether the videos are sourced from a broad range of open domains (Open-domain), and whether provide subtitle together with audio information (Sub.&Aud.). Video-MME-S/M/L denotes the short/medium/long part. It is important to note our comparison focuses solely on the multiple-choice subset of the datasets.
@@
The rapid development of MLLMs in recent years [66] has highlighted their impressive perception and cognitive capabilities across various multimodal benchmarks [16, 67, 43]. These advancements show the great potential of MLLMs to serve as a foundation that can digest the multimodal real world and pave the way toward artificial general intelligence. However, current MLLMs and their evaluation primarily focus on static visual data understanding, which fails to capture the dynamic nature of the real world involving complex interactions between objects over time. To approximate real-world scenarios more accurately, it is crucial to explore and assess the capabilities of MLLMs on sequential visual data, such as videos. Many early efforts [55, 33, 18] have been made to inspire the video understanding potentials of MLLMs with promising results. However, existing video-based benchmarks [32, 27, 45, 41] are still limited to thoroughly reveal their performance, such as a lack of diversity in video types, insufficient coverage of temporal dynamics, and the narrow focus on a single modality. These inevitably hinder the all-around evaluation of MLLMs.
@@
To this end, we introduce Video-MME, the first-ever comprehensive Multi-Modal Evaluation benchmark crafted for MLLMs in Video analysis. As exemplified in Figure 1, we meticulously curate a dataset of 900 videos across various scenarios, and annotate a set of 2,700 high-quality multiple-choice questions (3 per video) to foster a robust evaluation. As presented in Figure 2, for generalizability, our dataset widely spans 6 visual domains, including Knowledge, Film & Television, Sports Competition, Artistic Performance, Life Record, and Multilingual, with 30 fine-grained categories, e.g., technology, documentary, news report, esports, magic show, and fashion. Importantly, the videos vary significantly in length, ranging from 11 seconds to 1 hour, specifically evaluating the adaptability of MLLMs across varying temporal contexts. Furthermore, Video-MME enriches the assessment by incorporating the associated subtitles and audio tracks, thereby enhancing the analysis of multimodal inputs for video understanding.
@@
Using Video-MME, we benchmark various state-of-the-art MLLMs, including GPT-4V [48], GPT-4o [49], and Gemini 1.5 Pro [55], alongside open-source image models like InternVL-Chat-V1.5 [11] and video models like LLaVA-NeXT-Video [73]. Our experiments in Table 4 indicate that Gemini 1.5 Pro is the highest-performing commercial model, achieving an average accuracy of 75%. In comparison, open-source MLLMs exhibit substantial gaps compared to commercial models. For instance, the leading open-source model, VILA-1.5 [34], attains an overall accuracy of 59%. These findings suggest there is considerable room for improvement in the open-source community. Our benchmark is also available to advanced image-based models by extending their input to multi-frame images, e.g., Qwen-VL-Max [5] and InternVL-Chat-V1.5 [11]. The accuracies of both the models reach 50% , which is close to that of the video-specific model LLaVA-NeXT-Video, indicating that image understanding is the basis of video understanding and the wide applicability of Video-MME in the field of MLLMs. Further observations in Table 5 indicate that integrating subtitles and audios significantly enhances video comprehension capabilities, e.g., boosting Gemini 1.5 Pro by 6.2% and 4.3% respectively, with the gains being more pronounced for longer videos. A fine-grained analysis of task types reveals that subtitles and audios are particularly beneficial for videos requiring substantial domain knowledge. We also note a general decline in MLLM performance with increasing video length. This trend suggests that limitations in processing longer video sequences could be a critical bottleneck in the performance of MLLMs.
@@
2 Related Work
Advancements in MLLMs.
Recent advancements in MLLMs have seen notable progress [66, 19, 22, 59, 7]. MLLMs typically comprise three core modules: (i) a vision encoder for visual feature extraction, (ii) a modality alignment module to integrate visual features into the embedding space of the language model, and (iii) an LLM backbone for decoding multi-modal context. CLIP [51] and SigLIP [70] are widely-used for image encoding, while LLaMA [56] and Vicuna [12] serve as popular choices for LLMs. The alignment module varies from simple linear projections [35] to more complex architectures such as Q-Former [25, 13], and gated cross-attention layers substantiated by Flamingo [1, 4]. Additionally, Fuyu-8B [6] introduces a novel framework mapping raw image pixels directly to the LLM embedding space. Regarding MLLMs for processing videos [26, 44, 54], the key difference lies in how they encode the video into vision tokens compatible with the LLMs. Representative work like Video-LLaMA [71] first uses a ViT [14] with an image Q-Former to encode individual frames and then employs a video Q-Former for temporal modeling. VideoChat2 [27] utilizes a video transformer to encode video features and subsequently implements a Q-Former [25] to compress video tokens. To empower video MLLMs with temporal localization capability [20, 50, 60], TimeChat [53] constructs time-sensitive instruction tuning datasets and encodes timestamp knowledge into visual tokens. However, the potential of MLLMs in processing sequential visual data is still under-explored.
MLLM Benchmarks.
Alongside advancements in architecture, significant efforts have been made to improve benchmarking for MLLMs, guiding the development of the next generation of these models. Previous studies have integrated various aspects of evaluation, such as perception and cognitive capabilities, to create comprehensive benchmarks for assessing image MLLMs [16, 67, 40]. As image MLLMs have demonstrated exceptional performance in general perception tasks, benchmarks regarding scientific understanding [31], multi-modal mathematical reasoning [43, 72], and multi-disciplinary [69] capabilities have drawn increasing attention. For video MLLMs, similar efforts have been made to incorporate existing benchmarks [58, 29] for evaluating video understanding [27, 47]. Due to the inherently temporal characteristics of video modalities, specialized benchmarks have been created to evaluate temporal comprehension, underscoring the current limitations of video MLLMs in understanding video content [32, 41, 62, 57, 15, 3].
@@ 3 Video-MME
Figure 2: Statistics analysis of Video-MME. (left) Video categories. Our benchmark covers 6 key domains and 30 sub-class video types. (right) Video duration and question type distributions. Video-MME has a full spectrum of video length and covers different core abilities of MLLMs, enabling a comprehensive evaluation of temporal understanding.
Dataset Question Options Answer Subtitles
Short 11.5 17.2 4.0 198.6
Medium 12.2 20.6 5.0 1425.6
Long 14.5 31.0 7.5 6515.6
All 12.7 22.9 5.5 3086.5
Table 2: Average word counts of different textual fields in Video-MME. The word counts for questions, options, and answers are consistent across video lengths, indicating a standardized QA format and an increasing trend of subtitles information volume.
@@
3.1 Dataset Construction
The construction of the Video-MME dataset involves three main steps: video collection, question-answer annotation, and quality review. The details are as follows.
Video Collection.
To ensure comprehensive coverage across diverse video types, we first establish a domain hierarchy for sourcing raw videos from YouTube. We identify 6 key domains: Knowledge, Film & Television, Sports Competition, Life Record, and Multilingual, informed by popular trends on YouTube. Each domain is further subdivided into specific tags, such as football and basketball under Sports Competition, yielding a total of 30 finely-grained video categories. The complete domain-tag hierarchy and its distribution are illustrated on the left side of Figure 2. For each category, we collect videos with varying duration lengths, including short (< 2 minutes), medium (4-15 minutes), and long videos (30-60 minutes). Besides, we also obtain corresponding meta-information such as subtitles (if provided) and audios for in-depth investigation. The resulting dataset consists of 900 videos, including 744 videos with subtitles and all 900 with audios, representing a range of domains with a relatively balanced distribution of video durations, as depicted on the right side of Figure 2.
Question-Answer Annotation.
Following the collection of raw video data, we annotate the videos with high-quality question-answer (QA) pairs to rigorously assess the proficiency of MLLMs in video content interpretation. We employ a multiple-choice QA format to facilitate a straightforward and flexible assessment and recruit annotators with strong English proficiency and extensive research experience in vision-language learning. Specifically, each annotator is tasked with thoroughly viewing the entire video, then iteratively creating three relevant questions, each accompanied by four candidate options, contributing to a total of 2,700 QA pairs. As shown in the bottom right corner of Figure 2, these questions encompass 12 distinct task types, including perception, reasoning, and information synthesis. Each QA pair is designed to be closely tied to the video content, preventing MLLMs from answering accurately without directly referencing the video.
Quality Review.
To guarantee the quality of our dataset, we conduct a rigorous manual review process. First, each QA pair undergoes examination by a different annotator to verify that (i) the language expression is clear and precise; (ii) the question is answerable with logically consistent candidate options, and the designated correct answer is appropriate. Furthermore, to ensure that the questions are sufficiently challenging enough and require video content as a critical clue to answer [72], we provide the text-only questions to Gemini 1.5 Pro and filter out QA pairs that can be answered solely based on the textual questions. For instance, a question like “What is the biggest achievement of Argentina’s number 10 in 2022?” which can be inferred as the World Cup victory, would be excluded in this phase. Questions that do not meet this criterion are returned to the original annotator for revision. Statistical analysis shows that Gemini 1.5 Pro achieves less than 15% accuracy in the text-only setup, underscoring the robustness of the video content-based requirement. Through our dataset construction process, we strive to deliver a high-quality, diverse, and well-balanced dataset that will serve as a valuable resource for advancing research in multimodal understanding.
3.2 Dataset Statistics
In this section, we present the detailed statistics of our dataset to provide a more comprehensive understanding, including the meta information, QA pairs, certificate lengths, qualitative analysis, and comparison to previous works.
Video & Meta Information.
Our dataset comprises a total of 900 videos, 744 subtitles, and 900 audio files. Most videos are accompanied by both subtitles and audios, providing valuable resources for investigating the impact of external information on video understanding performance. The upper right part of Figure 2 illustrates the duration distribution of the collected videos. Specifically, within the short video category, longer videos occupy a larger portion. For medium-length videos, the duration distribution is more uniform, while long videos exhibit a long-tailed distribution, with fewer samples as the duration increases. The bottom right part of Figure 2 shows the distribution of task types. Shorter videos predominantly involve perception-related tasks such as action and object recognition, while longer videos mainly feature tasks related to reasoning. Overall, this analysis highlights that Video-MME covers a wide range of video durations and task types, enabling a comprehensive evaluation of temporal understanding.
QA Pairs.
We present a detailed analysis of language diversity in the questions and answers within our dataset. Table 2 lists the average word counts of the textual fields including questions, options and answers in our dataset. Notably, the word counts display notable consistency across different video lengths, suggesting a standardized format in the QA pairs. In contrast, subtitle word counts increase markedly with video length—e.g., short videos average 198.6 words, while long videos reach up to 6.5K words. This trend indicates longer videos contain more information, as evidenced by the increased volume of subtitles. The analysis reveals that our questions are diverse, and the answers are well-balanced. In addition, the distribution of the options (A/B/C/D) follows a near-uniform distribution (25.1/27.2/25.3/22.4%), ensuring an unbiased evaluation.
@@ 3.2 Dataset Statistics
Inspired by the recent work EgoSchema [45], we adopt the certificate length to analyze the temporal difficulty of the QA pairs. The certificate of a given video QA pair is defined as the minimum set of sub-clips of the video that are both necessary and sufficient to convince a human verifier that the marked annotation is correct. The certificate length is calculated as the sum of the temporal lengths of the sub-clips identified. We randomly sample 3 videos from each class and calculate the certificate length distribution. As shown in Table 3, our dataset yields a median certificate length of 26s, 164.7s, and 890.7s for short, medium and long videos, respectively. Compared with the certificate length of EgoSchema, our medium and long video subset requires much longer video content digestion to answer the question. To the best of our knowledge, this analysis makes our Video-MME the most challenging Video QA dataset to date.
[…]
Table 4: Performance of MLLMs on Video-MME. The results are evaluated on short, medium, and long durations, under the settings of “without subtitles” and “with subtitles”. In the “with subtitles” setting, subtitles are selected to correspond to the sampled video frames.